OpenStudy (anonymous):

Find the coordinates of two points on the line with the given equation. Then use the points to find the slope of the line. Questions: 1)y= -3 + 11 2) y = -17 3) y = 7/8x - 11 4) 9x + 8y = 56 5) x=10 6) 7y - 3x = -147 >Thank you in advance

OpenStudy (anonymous):

I won't tell you the answers, but I will tell you how to solve it. Do you have a graphing calculator? If so, go to "y=", type in the equation after the equal sign, and then press"2nd" then "graph". If not, use this site: http://www.meta-calculator.com/online/ Select "graphing calculator", type in the numbers after equals in the first "y=" box. Select "table" Scroll down and seclect "create table". Those are your points. To find the slope, you take the two points and use this formula: \[\frac{ y ^{2}-y ^{1} }{ x ^{2}-x ^{1} }\] Example (#1): 1)y= -3 + 11 2 points are (-7,8) and (-6,8) Now for the slope formula, you will subtract the the first y point from the second and put it over the second x point minus the first x point. So: \[\frac{ 8-8 }{ -6+7 }\] (since two negatives make a positive, - -7 would be +7). So your answer is \[\frac{ 0 }{ 1 }\] And since zero is on top, your slope is zero. If zero was on bottom, your slope would be undefined.